Wiater Company operates a small manufacturing facility. On January 1, 2018, an asset account for the company showed the folloJournal entry worksheet < A Record the adjusting entry for depreciation expense at year-end 2017. Note: Enter debits before cRequired 1 Required 2 Required 3 Starting at the beginning of 2018, what is the remaining estimated life? Remaining EstimatedJournal entry worksheet < A B Record the expenditure for ordinary repairs incurred. Note: Enter debits before credits. TransaJournal entry worksheet Record the expenditure for extraordinary repairs incurred. Note: Enter debits before credits. Transac
Answer #1
1
Debit Credit
Depreciation expense 12400 =(200000-14000)/15
       Accumulated Depreciation-Equipment 12400
2
Years of use 5 =62000/12400
Remaining estimated life 10 years =15-5
3
Debit Credit
Repairs and Maintenance expense 2450
        Cash 2450
Equipment 27000
        Cash 27000